# A fast particle-field method for 3D Chemotaxis systemsJingyuan Hu, Zhongjian Wang, Jack Xin and Zhiwen Zhang

### v2

@@ 5.2 Critical Mass in 2D and 3D Systems
Recall the two-dimensional parabolic-elliptic Keller-Segel system with parameters $(k,\epsilon,\mu,\chi)=(0,0,1,1)$, where the critical mass threshold $M_{\text{crit}}=8\pi$ governs finite-time blow-up (see Section 2.1). Numerical simulations initialize with uniform distributions on the unit disk $\rho_{0}(\mathbf{x})=\frac{M_{0}}{\pi}\mathbf{1}_{|\mathbf{x}|\leq 1}$, for masses $M_{0}\in\{22.0,23.0,\ldots,27.0\}$, discretized in the spatial domain $\Omega=[-4,4]^{2}$ with grid resolution $H=2048$ on each axis and $P=2^{20}$ particles, utilizing fourth-order Particle-to-Grid and second-order Grid-to-Particle interpolations, coupled with time integration $\tau=5\times 10^{-4}$ up to final time $T=2.0$, totaling $N_{T}=4000$ steps.
@@
Figure 5: Two-dimensional particle distributions at $T=2.0$ for different initial masses. (a)-(d): Subcritical to near-critical regimes ($M_{0}=22.0$$25.0$); (e)-(f): Supercritical regimes ($M_{0}=26.0$$27.0$).
@@
For the three-dimensional system (1) with parameters
$(k,\epsilon,\mu,\chi)=(10^{-1},10^{-4},1,1),$
no explicit critical mass threshold is theoretically established. Numerical experiments are conducted with initial masses $M_{0}\in\{45,46,\ldots,50\}$ uniformly distributed within the unit ball $\rho_{0}(\mathbf{x})=\frac{3M_{0}}{4\pi}\mathbf{1}_{\|\mathbf{x}\|\leq 1}$. The spatial domain $\Omega=[-4,4]^{3}$ is discretized using $H=256$ grid points per axis and $P=2^{20}$ Lagrangian particles.
The computational framework combines fourth-order Particle-to-Grid and second-order Grid-to-Particle interpolations, with time integration performed at fixed step $\tau=5\times 10^{-4}$ up to final time $T=1.0$, totaling $N_{T}=2000$ steps. The particle distributions at the final time corresponding to different initial masses are shown in Figure 7.
Figure 7 indicates a critical mass threshold between $M_{0}=48.0$ and $M_{0}=49.0$, consistent with the results of finite-difference solutions of the spherically symmetric 1D problem.
Figure 7: Three-dimensional particle distributions at $T=1.0$ for different initial masses. (a)-(d): Subcritical to near-critical regimes ($M_{0}=45.0$$48.0$); (e)-(f): Supercritical regimes ($M_{0}=49.0$$50.0$).
The original SIPF method [SIPF1] demonstrates that while blow-up occurs, the computed $\|\hat{c}\|_{\infty}$ grows with the truncation frequency $H_{0}$. Figure 8 shows this indicator for 2D and 3D cases, confirming critical masses consistent with particle distributions in Figures 5 and 7.
Figure 8: $\frac{\|\hat{c}\|_{\infty,H_{0}=32}}{\|\hat{c}\|_{\infty,H_{0}=8}}$ vs. computation time $T$.
@@ 5.3 Blowup Dynamics in Tetrahedral Configuration
preserving tetrahedral symmetry with edge length $\sqrt{3}$. The parameters of the system (1) remain $\mu=\chi=1$, $\epsilon=10^{-4}$, $k=10^{-1}$ as in Section 5.1.
Figure 9: Initial tetrahedral density configuration
We tested $M_{0}\in\{20,28,40,56,80,112,160\}$ with total simulation time $T=0.4$ and timestep $\tau=10^{-4}$ ($N_{T}=4000$). The grid resolution is fixed at $H=256$. Due to the near blowup scenarios, the smoothness of $\nabla c$ is relatively poor. Therefore, we utilize fourth-order Particle-to-Grid interpolations and second-order Grid-to-Particle interpolations to avoid singularities. Particle configurations at $T=0.1$, $0.2$, and $0.4$ are shown in Fig. 10 for three regimes: $M_{0}=20$ (non-blowup), 80 (central blowup), and 160 (two-stage blowup), with other values demonstrating analogous behavior within their respective categories.
Three distinct dynamical regimes emerge. For small $M_{0}$, no blowup occurs. At intermediate $M_{0}$, the four clusters merge into a single aggregate prior to a unified blowup. For large $M_{0}$, the system exhibits a two-stage process: simultaneous blowups initially develop at the original tetrahedral positions, followed by coalescence into a central singularity.
@@
Figure 10: Blowup patterns for different $M_{0}$
@@ 5.4 Collapsing Ring Blowup Dynamics
At the final stage, we observe a special blow-up pattern where the mass concentrates on an evolving ring $x_{1}^{2}+x_{2}^{2}=r(t)^{2}$, with the radius $r(t)$ decaying as $t\to T$.
Numerical simulations initialize with a uniform distribution inside the toroidal volume defined by $(1.0-\sqrt{x_{1}^{2}+x_{2}^{2}})^{2}+x_{3}^{2}\leq 0.4^{2}$, with parameters $(k,\epsilon,\mu,\chi)=(10^{-1},10^{-4},1,1)$ in system (1), initial mass $M_{0}=180$, and final time $T=6\times 10^{-2}$. The spatial domain $\Omega=[-4,4]^{3}$ is discretized with $H=256$ grid points per axis, using $P=2^{20}$ particles, fourth-order Particle-to-Grid and second-order Grid-to-Particle interpolations, with time step $\tau=0.5\times 10^{-4}$, equivalently $N_{T}=1200$.
Figure 11 shows the projected scatter plot of the particle distributions in the $x_{1}$-$x_{2}$ plane in different time steps. The mass progressively collapses into a thin ring with linear density as $t\to T$. After complete collapse, numerical stability degrades severely, revealing extreme sensitivity to non-symmetric perturbations. This instability drives particles to aggregate into point-like clusters (singularities) within localized regions, each with a radius of approximately one grid length. High particle counts and grid resolution are necessary to observe the ring structure. The original SIPF method ($P=10^{4},H=32$) is not sufficient to observe the ring structure, as the particles undergo a blow-up at the origin before aggregating into a ring.
Figure 11: Toroidal collapse evolution
@@ 6 Conclusions
In this work, we introduced the SIPF-PIC framework, a hybrid particle/spectral/particle-in-cell method that synergizes localized particle-grid interpolations with global spectral field solvers accelerated by FFT. This approach achieves a computational complexity of $O(P+H^{3}\log H)$, enabling efficient simulations of 3D Keller-Segel systems with large particle numbers and fine grids. A unified error analysis framework was developed to rigorously quantify cumulative approximation errors arising from particle-grid interpolations and spectral field evaluations, establishing convergence guarantees for non-blow-up cases.
The SIPF-PIC method successfully resolves complex blow-up dynamics beyond classical single-point singularities, such as ring-shaped blow-up patterns in which mass is concentrated on an evolving ring structure. Numerical experiments, including convergence tests and 3D visualizations, validate the enhanced precision and scalability of the method to capture multiscale phenomena.
Future work will focus on extending the SIPF-PIC framework to other related systems with collective particle-field interactions. Furthermore, our objective is to develop adaptive strategies to improve the robustness of the method in handling near-blow-up cases, where high-resolution tracking of density singularities remains a critical challenge.
@@ 7 Acknowledgements
ZW was partially supported by NTU SUG-023162-00001 and MOE AcRF Tier 1 Grant RG17/24. JX was partially supported by NSF grant DMS-2309520. ZZ was partially supported by the National Natural Science Foundation of China (Projects 92470103 and 12171406), the Hong Kong RGC grant (Projects 17304324 and 17300325), the Seed Funding Programme for Basic Research (HKU), and the Hong Kong RGC Research Fellow Scheme 2025. The computations were performed at the research computing facilities provided by Information Technology Services, the University of Hong Kong.

### v3

@@ 5.2 Critical mass in 2D and 3D systems
Recall the 2D parabolic-elliptic KS system with parameters $(k,\epsilon,\mu,\chi)=(0,0,1,1)$, where the critical mass threshold $M_{\text{crit}}=8\pi$ governs finite-time blow-up (see Section 2.1). Numerical simulations initialize with uniform distributions on the unit disk $\rho_{0}(\mathbf{x})=\frac{M_{0}}{\pi}\mathbf{1}_{|\mathbf{x}|\leq 1}$, for masses $M_{0}\in\{22.0,23.0,\ldots,27.0\}$, discretized in the spatial domain $\Omega=[-4,4]^{2}$ with grid resolution $H=2048$ on each axis and $P=2^{20}$ particles, utilizing fourth-order particle-to-grid and second-order grid-to-particle interpolations, coupled with a time step $\tau=5\times 10^{-4}$ up to final time $T=2.0$, totaling $N_{T}=4000$ steps.
@@
Figure 5: Two-dimensional particle distributions at $T=2.0$ for different initial masses. (a)-(d): Subcritical to near-critical regimes ($M_{0}=22.0$-$25.0$); (e)-(f): Supercritical regimes ($M_{0}=26.0$-$27.0$).
@@
For the 3D system (1) with parameters $(k,\epsilon,\mu,\chi)=(10^{-1},10^{-4},1,1)$, no explicit critical mass threshold is theoretically established. Numerical experiments are conducted with initial masses $M_{0}\in\{45,46,\ldots,50\}$ uniformly distributed within the unit ball $\rho_{0}(\mathbf{x})=\frac{3M_{0}}{4\pi}\mathbf{1}_{\|\mathbf{x}\|\leq 1}$. The spatial domain $\Omega=[-4,4]^{3}$ is discretized using $H=256$ grid points per axis and $P=2^{20}$ particles.
The computational framework combines fourth-order particle-to-grid and second-order grid-to-particle interpolations, with time integration performed at a fixed time step $\tau=5\times 10^{-4}$ up to final time $T=1.0$, totaling $N_{T}=2000$ steps. The particle distributions at the final time corresponding to different initial masses are shown in Figure 7.
Figure 7 indicates a critical mass threshold between $M_{0}=48.0$ and $M_{0}=49.0$, consistent with the results of finite-difference solutions to the spherically symmetric 1D problem.
Figure 7: Three-dimensional particle distributions at $T=1.0$ for different initial masses. (a)-(d): Subcritical to near-critical regimes ($M_{0}=45.0$-$48.0$); (e)-(f): Supercritical regimes ($M_{0}=49.0$-$50.0$).
The original SIPF method [SIPF1] demonstrates that while blow-up occurs, the computed $\|\hat{c}\|_{\infty}$ grows with the truncation frequency $H_{0}$. Figure 8 shows this indicator for 2D and 3D cases, confirming critical masses consistent with those derived from particle distributions in Figures 5 and 7.
Figure 8: $\frac{\|\hat{c}\|_{\infty,H_{0}=32}}{\|\hat{c}\|_{\infty,H_{0}=8}}$ vs. computational time $T$.
@@ 5.3 Blowup dynamics in tetrahedral configuration
preserving tetrahedral symmetry with edge length $\sqrt{3}$. The parameters of the system (1) remain $\mu=\chi=1$, $\epsilon=10^{-4}$, and $k=10^{-1}$ as in Section 5.1.
Figure 9: Initial tetrahedral density configuration.
We performed simulations for $M_{0}\in\{20,28,40,56,80,112,160\}$ with total simulation time $T=0.4$ and a time step $\tau=10^{-4}$ ($N_{T}=4000$). The grid resolution is fixed at $H=256$. Due to the near blowup scenarios, the smoothness of $\nabla c$ is relatively poor. Therefore, we utilize fourth-order particle-to-grid interpolations and second-order grid-to-particle interpolations to avoid spurious singularities. Particle configurations at $T=0.1$, $0.2$, and $0.4$ are shown in Fig. 10 for three regimes: $M_{0}=20$ (non-blowup), 80 (central blowup), and 160 (two-stage blowup), with other values demonstrating analogous behavior within their respective categories.
Three distinct dynamical regimes are observed. For small $M_{0}$, no blowup phenomenon occurs. At intermediate $M_{0}$, the four clusters merge into a single aggregate prior to a single unified blowup at the center. For large $M_{0}$, the system exhibits a two-stage process: simultaneous blowups initially develop at the original tetrahedral positions, followed by coalescence into a central singularity.
@@
Figure 10: Blowup patterns for different $M_{0}$.
@@ 5.4 Collapsing ring blowup dynamics
In the last experiment, we investigate a special blow-up pattern where the mass concentrates on an evolving ring $x_{1}^{2}+x_{2}^{2}=r(t)^{2}$, with the radius $r(t)$ decaying as $t\to T$.
Numerical simulations initialize with a uniform distribution inside the toroidal volume defined by $(1.0-\sqrt{x_{1}^{2}+x_{2}^{2}})^{2}+x_{3}^{2}\leq 0.4^{2}$, with parameters $(k,\epsilon,\mu,\chi)=(10^{-1},10^{-4},1,1)$ in system (1), an initial mass $M_{0}=180$, and a final computational time $T=6\times 10^{-2}$. The spatial domain $\Omega=[-4,4]^{3}$ is discretized with $H=256$ grid points per axis, using $P=2^{20}$ particles, fourth-order particle-to-grid and second-order grid-to-particle interpolations, with a time step $\tau=0.5\times 10^{-4}$, equivalently $N_{T}=1200$.
Figure 11 shows the projected scatter plot of the particle distributions in the $x_{1}$-$x_{2}$ plane at different time steps. The mass progressively collapses into a thin ring with an asymptotically linear density profile as $t\to T$. After complete collapse, numerical stability degrades severely, revealing an extreme sensitivity to asymmetric perturbations. This instability drives particles to aggregate into point-like clusters (singularities) within localized regions, each with a radius of approximately one grid spacing. High particle number and grid resolution are necessary to resolve the ring structure. The original SIPF method ($P=10^{4},H=32$) is not sufficient to resolve the ring structure, as the particles undergo a blow-up at the origin rather than aggregating into a ring.
Figure 11: Toroidal collapse evolution.
@@ 6 Conclusions
In this paper, we develop the SIPF-PIC method, a hybrid particle-spectral particle-in-cell method that combines localized particle-grid interpolations with global spectral field solvers accelerated by the FFT method. This new method achieves a computational complexity of $O(P+H^{3}\log H)$, enabling efficient simulations of 3D Keller-Segel (KS) systems with large particle numbers and fine grids. A unified error analysis framework is established to rigorously quantify approximation errors arising from particle-grid interpolations and spectral field evaluations, thus providing convergence guarantees for our SIPF-PIC method in computing non-blow-up solutions of parabolic-parabolic KS systems.
The SIPF-PIC method successfully resolves complex blow-up dynamics beyond classical single-point singularities, such as ring-shaped blow-up patterns in which mass concentrates on an evolving ring structure. Numerical experiments, including convergence tests and 3D visualizations, demonstrate the enhanced precision and scalability of the method in capturing aggregation and near-singular phenomena.
Future work will focus on extending the SIPF-PIC method to other related systems with collective particle-field interactions. Furthermore, we aim to develop adaptive strategies to improve the robustness of the method in handling near-blow-up cases, where high-resolution tracking of density singularities remains a critical challenge.
@@ 7 Acknowledgements
ZW was partially supported by NTU SUG-023162-00001 and MOE AcRF Tier 1 Grant RG17/24. JX was partially supported by NSF Grant DMS-2309520, the Swedish Research Council Grant No. 2021-06594 at the Institut Mittag-Leffler in Djursholm, Sweden, and the E. Schrödinger Institute in Vienna, Austria, both during his stay in the fall of 2025. ZZ was partially supported by the National Natural Science Foundation of China (Projects 92470103 and 12171406), the Hong Kong RGC Grant (Projects 17304324 and 17300325), the Seed Funding Programme for Basic Research (HKU), and the Hong Kong RGC Research Fellow Scheme 2025. The computations were performed at the research computing facilities provided by the Greenplanet Cluster at UC Irvine and the Information Technology Services, The University of Hong Kong.